OpenStudy (masonsurfs):

Use the quadratic formula to solve the equation -4x^2-3x+2=0 Please explain if you can Thanks

OpenStudy (wolf1728):

-4x^2-3x+2=0 a=-4 b=-3 c=2 x=[-b +-sqrt(b^2 - 4*a*c)] / 2a x= [--3 +-sqrt(-3^2 -(4*-4*2))]/2*-4 x = [3 +-sqrt(9 -(-32)] / -8 x = [3 +- sqrt(41)]/ -8

OpenStudy (wolf1728):

x = (3 + sqrt(41))/ -8 x = (3 - sqrt(41))/ -8

OpenStudy (wolf1728):

Does that help Mason?

OpenStudy (masonsurfs):

Yes kind of thank you for explaining

OpenStudy (wolf1728):

u r welcome